a new truck can be paid for by 48 monthly payments of x dollars each plus a down payment of 2.5 times the amount of the monthly payment. This implies that the total amount paid for the truck is 2.5x + 48x.
a. Use the distributive property to rewrite the expression.
b. What is the total amount paid for a truck that has a monthly payment of $435?

Answer by rapaljer(4671)

2.5x + 48x
x(2.5 + 48)
x(50.5)
50.5x

(50.5)*(435) = $21967.50

Check:
Down Payment = 2.5*435 = $1087.50
Payments = 48* 435 = $20880
Total = $21967.50

R^2 at SCC